OpenStudy (anonymous):

Solve: −1/2 x + 2 = −x + 7

sam (.sam.):

\[\begin{array}{l} 2-\frac{x}{2}=7-x \\ \text{Add }(x-2)\text{ to both sides} \\ \frac{x}{2}=5 \\ \text{Divide both sides by }\frac{1}{2} \\ x=10 \\\end{array}\]

OpenStudy (mayankdevnani):